Two moles of nitrogen are in a 3-liter container at a pressure of 5.0Í106 Pa. Find the average translational kinetic energy of a molecule.


1
Expert's answer
2021-06-03T09:18:12-0400
"E=\frac{3}{2}kT\\E=\frac{3pV}{2N}\\E=\frac{3pV}{2\nu N_A}
\\E=\frac{3(3\cdot10^{-3})(5\cdot10^6)}{2(2)(6.04\cdot10^{23})}=1.9\cdot10^{-20}\ J"